solve
8 9 1
-- = -- - -- so it reads 8 over y equals 9 over y minus 1 over 5.
y y 5

Answer by checkley77(12844)

8 9 1
-- = -- - --
y y 5
8/Y=9/Y-1/5
8/Y-9/Y=-1/5
-1/Y=-1/5 CROSS MULTIPLY.
-Y=-5
Y=5 ANS.
PROOF:
8/5=9/5-1/5
8/5=8/5